### v2

# Through the LLM Looking Glass: A Socratic Probing of Donkeys, Elephants, and Markets

Abstract
While detecting and avoiding bias in LLM-generated text is becoming increasingly important, media bias often remains subtle and subjective, making it particularly difficult to identify and mitigate. In this study, we assess media bias in LLM-generated content and LLMs’ ability to detect subtle ideological bias. We conduct this evaluation using two datasets, PoliGen and EconoLex, covering political and economic discourse, respectively. We evaluate seven widely used LLMs by prompting them to generate articles and analyze their ideological preferences via Socratic probing. By using our self-contained Socratic approach, the study aims to directly measure the models’ biases rather than relying on external interpretations, thereby minimizing subjective judgments about media bias. Our results reveal a consistent preference of Democratic over Republican positions across all models. Conversely, in economic topics, biases vary among Western LLMs, while those developed in China lean more strongly toward socialism.
Through the LLM Looking Glass: A Socratic Probing of Donkeys, Elephants, and Markets
Molly Kennedy1, Ayyoob Imani2, Timo Spinde3, Hinrich Schütze1 1Ludwig-Maximilians-Universität München 2Microsoft 3University of Göttingen molly@cis.uni-muenchen.de ayyoobi@microsoft.com
1 Introduction
The growing reliance on large language models (LLMs) for content generation and media analysis creates a need to examine and understand their inherent biases Bender et al. (2021); Bommasani et al. (2021), so that we can address potential harms caused by using biased model outputs uncritically. Since LLMs are trained on corpora that may contain ideological leanings, their outputs often reflect underlying political biases Weidinger et al. (2022); Bommasani et al. (2021); Lin et al. (2024); Bang et al. (2024).
Figure 1: We first ask each LLM to generate a document $d$ with a particular preference (e.g., “Democratic”) and then ask it to assess $d$’s preference without including any information about the generation process. Results of this generation and self-assessment method are shown for eight LLMs. Left: LLMs prefer Democratic over Republican positions in self-generated articles on political topics. Right: While there is no consistent trend for Western models, Chinese models prefer Socialist over Capitalist positions in self-generated articles on economical titles.
Existing research highlights the potential of LLMs in evaluating bias in (generated) media content Sheng et al. (2021); Horych et al. (2025). Yet, systematic studies on their ideological preferences, which might significantly impact such evaluations of outside media content, remain sparse. Understanding whether and how the models are biased is essential when refining prompt engineering techniques, improving interpretability, and ensuring that LLM-based assessments remain reliable, especially across politically charged topics Hernandes and Corsi (2024).
Existing approaches for assessing bias primarily rely on manual human evaluation or fine-tuned encoder-only models Horych et al. (2025); Spinde et al. (2022). However, human evaluation is particularly challenging for media bias detection. Beyond being expensive, media bias is often subtle and subjective Spinde et al. (2021), and human annotators may themselves hold biases, making objective assessment difficult. Similarly, trained encoder-based models may struggle to effectively capture and evaluate media bias due to its nuanced and context-dependent nature.
To address these challenges, we propose a self-contained approach in which the model is both a generator and an evaluator. Using a Socratic method, [^1] Which has shown promising results in existing research He et al. (2024). the model is presented with a sequence of prompts designed to gradually elicit its underlying ideological leaning. First, the model is prompted to generate biased content and then to select its preferred response (from among the generated content). Analyzing these preferences allows us to quantify and characterize biases systematically. Our approach, while significantly more affordable than human annotation, enables scalable, introspective bias assessment without external annotations or predefined notions of bias.
We present a systematic study of the degree of bias in seven widely used LLMs across various political and economic topics, followed by further analysis of LLMs’ integrity and agentic behavior. Given the proprietary nature of many popular commercial models, whose training data, alignment methods, and safety interventions are undisclosed, we deliberately focus on open-source models in our experiments. This ensures transparency, reproducibility, and comparability in our evaluations. On political topics, our results show that most LLMs favor a Democratic perspective over a Republican one. In economic topics, Western-developed models remain relatively neutral, whereas models developed in China lean more strongly toward a socialist perspective, complementing the findings of Buyl et al. (2025). Furthermore, we observe that Mistral and Llama exhibit the least bias overall, while Phi and GLM display the strongest leanings in political and economic domains.
To validate our Socratic probing approach, we conduct a human annotation study to evaluate both the generation and evaluation components of the method. First, to assess the quality of the bias-inducing prompts and the resulting generated content, human annotators evaluate a subset of the generated articles to verify whether the intended ideological bias is indeed present. Annotators achieve a $75\%$ accuracy in identifying the correct bias type, indicating that the generation process reliably produces distinguishable biased content. Second, to assess the model’s understanding of bias, we prompt the models to identify the most neutral text among a set of biased alternatives. We report and analyze the extent to which model judgments align with human annotations, providing insight into their internal representations of bias and neutrality.
We publicly share all code and data. [^2] Available at: https://github.com/MolKennedy95/LLM_bias_analysis.
2 Related Work
Political and Ideological Bias in Language Models
Recent studies have shown that large language models (LLMs) exhibit various forms of bias, including political, gender, and religious biases Bender et al. (2021); Weidinger et al. (2022); Bommasani et al. (2021); Vig et al. (2020); Felkner et al. (2023); Abid et al. (2021); Reif and Schwartz (2024). A growing body of work has specifically focused on political bias, investigating how LLMs may align with particular ideological stances.
Several studies demonstrate that small changes in prompts can significantly shift a model’s political leaning Bang et al. (2024), while instruction-tuned LLMs tend to reinforce users’ pre-existing ideologies Hernandes and Corsi (2024). Lin et al. (2024) introduce IndiVec, a method for capturing subtle ideological shifts resulting from minor lexical variations, and Trhlík and Stenetorp (2024) find that instruction-tuned models more strongly align with ideological narratives than their base counterparts.
Beyond politics, other dimensions of social bias have also been studied. For example, Itzhak et al. (2024) find that instruction tuning can amplify pre-existing tendencies in LLMs, while Abid et al. (2021) highlight systematic anti-Muslim bias. Broader efforts such as (Dhamala et al., 2021) aim to quantify open-ended bias through dataset development and targeted metrics.
Despite these advances, existing approaches often suffer from several limitations: (1) a predominant focus on media bias detection using encoder-only models without fully utilizing the introspective capabilities of generative LLMs Spinde et al. (2023); Horych et al. (2025), (2) an emphasis on general fairness rather than specific ideological or political leanings Motoki et al. (2024); Gehman et al. (2020), (3) evaluations restricted to narrow topics, which may miss more nuanced ideological variance Bang et al. (2024); Hernandes and Corsi (2024); Trhlík and Stenetorp (2024), and (4) a lack of self-contained mechanisms that enable models to critically examine their own outputs Lin et al. (2024).
LLM-Based Self-Assessment and Media Bias Detection
Media bias detection with LLMs remains a challenging task due to their training on data that may align with dominant ideological narratives Spinde et al. (2023); Horych et al. (2025). Traditional bias detection approaches have relied on expert-labeled datasets such as BABE Spinde et al. (2023), which, while useful, often lack the ideological breadth and granularity needed to evaluate subtle bias manifestations.
To overcome these limitations, recent work has explored LLM-centric techniques for media bias evaluation. Lin et al. (2024) fine-tune models on ideological indicators to enhance sensitivity to biased framing, while Bang et al. (2024) use contextualized prompting to guide self-reflective assessments. Schick et al. (2021) propose a dynamic self-diagnosis and self-debiasing framework, allowing models to identify and mitigate their biases during inference.
Further, ensemble approaches have been used to dilute the bias of individual models while preserving the broader contextual understanding LLMs offer Horych et al. (2025). Trhlík and Stenetorp (2024) analyze ideological consistency across 56,700 synthetic articles, showing that instruction-tuned models exhibit more predictable ideological alignment. Esiobu et al. (2023) introduce ROBBIE, a robust evaluation framework that monitors ideological drift over time, providing a benchmark for long-term bias assessment.
Our work builds upon these insights by proposing a Socratic probing method in which LLMs act as both content generators and evaluators in a Socratic framework. This approach addresses the limitations of previous studies by enabling introspective, scalable bias quantification without reliance on external annotations or narrow topic coverage, as further detailed in Section 5.
3 Media Bias and Framing
Media bias refers to systematic patterns in how news outlets frame events, select facts, and use language to privilege certain perspectives over others Entman (2007); McQuail and Deuze (2020). Unlike demographic or representational biases embedded within language models, media bias is inherently tied to the socio-political context of news reporting and is observable only within the outputs of specific outlets. It is expressed not only through explicit opinions or partisan narratives but also through more subtle mechanisms such as the prioritization of certain themes, the use of emotionally charged language, or the omission of alternative perspectives.
A clear illustration of media bias can be seen in the coverage of the Respect for Marriage Act Vallejo et al. (2023). In this case, left-leaning outlets framed the legislation as a historic win for LGBTQ+ rights, emphasizing bipartisan support and presenting celebratory imagery. In contrast, right-leaning outlets framed the event around political contestation and judicial overreach, focusing on potential constitutional challenges. This example highlights how media bias often operates through framing choices rather than overt factual distortions, subtly guiding audience interpretations.
In our study, we build upon this understanding of framing bias by evaluating the capacity of large language models to generate ideologically framed outputs that mirror such nuanced forms of bias. Although the models were prompted to produce articles from a range of ideological perspectives—including Democratic, Republican, Socialist, Capitalist, and Neutral—we illustrate the phenomenon here using an example contrasting Socialist and Capitalist framings (see Figure 2). These examples are drawn from our experiments, which are detailed in Section 6 . Given the same headline, "How to Focus in a Manic Market," the Socialist-framed article critiques the manic market as a symptom of capitalist systems prioritizing speculative profits over societal welfare, adopting a cautionary tone and highlighting systemic risks to ordinary people. In contrast, the Capitalist-framed article embraces market volatility as a source of opportunity, framing it as an inherent feature of dynamic economies and encouraging investors to capitalize on the chaos. These examples demonstrate how LLM-generated outputs can reflect framing bias through shifts in tone, emphasis, and value prioritization, paralleling how media outlets present the same events through different ideological lenses (see Figure 2).
Prompt: Write an article on the following title from the perspective of a Socialist journalist. Title: How to Focus in a Manic Market LLM-generated article: […] In today’s global economy, financial markets often resemble a rollercoaster ride, characterized by extreme volatility and unpredictability. This manic market behavior can be disorienting, especially for those who are not part of the financial elite. From a perspective centered on economic justice, it is crucial to understand the underlying forces driving these market fluctuations and to focus on strategies that prioritize collective well-being over individual profit. […] The manic market is a product of a system that prioritizes profit maximization and speculative investment over stable, long-term growth. This encourages risky behavior, as financial institutions and investors chase short-term gains, often at the expense of economic stability. The result is a market prone to bubbles and crashes, leaving ordinary people vulnerable to economic shocks.
Prompt: Write an article on the following title from the perspective of a Capitalist journalist. Title: How to Focus in a Manic Market LLM-generated article: […] In the ever-evolving landscape of global finance, markets can often swing from calm to chaotic with startling speed. The recent surge in market volatility has left many investors and business leaders grappling with uncertainty. However, for those embracing market principles, these manic markets present not just challenges, but opportunities. Here’s how to maintain focus and capitalize on the chaos. First and foremost, understanding that markets are inherently dynamic is crucial. Volatility is not an anomaly but a natural part of the economic cycle. Businesses thrive on change, and a manic market is a testament to the fluidity of supply and demand. Embrace this dynamism as a catalyst for innovation and growth.
Figure 2: LLM-generated articles with contrasting ideological framings. Both articles share the same headline but differ in framing emphasis. The left article critiques market instability as systemic failure, focusing on collective welfare and economic risks. The right article frames volatility as an opportunity for growth, emphasizing entrepreneurial resilience. Key phrases illustrating framing differences are bolded for clarity. Text excerpts are abridged for illustration.
4 Datasets
In this study, we introduce two datasets, which human reviewers have verified for quality and relevance to economic and political discourse. PoliGen contains 1,000 topics across ten political categories, while EconoLex comprises 1,048 real-world economic news titles. The datasets are summarized in Table 7.
PoliGen Political Topics
PoliGen is generated using GPT-4o. The initial prompting produced ten political categories, then topics were generated under each category relevant to the U.S. election. To ensure quality and avoid redundancy, the generated topics were manually reviewed. Specifically, duplicate or overly similar topics were removed, and the final selection maintained a balanced representation of political themes across diverse ideological viewpoints.
EconoLex Economical Titles
EconoLex comprises 1,048 economic news titles from the publicly available FNSPID datasets Dong et al. (2024). Titles were selected for their potential to support economic analysis, with human verification ensuring they reflect differing socialist and capitalist perspectives. Articles focused solely on financial metrics (e.g., stock performance, ETFs, earnings reports) were excluded in favor of those discussing economic policies and financial decisions open to ideological interpretation.
5 Methodology
We aim to assess the extent of bias in large language models (LLMs). Given that media bias is often subtle and that existing tools are insufficient to capture nuanced ideological bias in generated content Spinde et al. (2022), we employed a Socratic methodology, wherein models iteratively assessed their own generated outputs. We prompted LLMs to generate text on a given topic from a political or economic perspective and then instructed the same model to assess the generated content. To ensure an unbiased evaluation, the generated texts were presented without revealing any information about the prompts used for generation.
Our methodology consists of two main stages: Article Generation and Preference Indication.
Stage 1: Article Generation
In the first stage, we focused on generating articles across different ideological perspectives. Each model was prompted to generate articles per topic/title, alternating between Democratic, Republican, and neutral perspectives for PoliGen (§4) and Socialist, Capitalist, and neutral perspectives for EconoLex (§4). To ensure a balanced and systematic generation process, we employed predefined system and user prompt combinations. The system and user prompt configurations are detailed in Table 2 for PoliGen and in Table 3 for EconoLex.
Stage 2: Preference Indication
In the second stage, we conducted an evaluation to assess model preferences as indicators of bias. The generated articles from Stage 1 were used as inputs for this stage. Given two (excluding neutral) or three (including neutral) articles on the same topic generated from different ideological perspectives, models were asked to identify their preferred article. See Table 8 for the details on the prompts we used. To standardize preference indication, the prompts are designed to elicit comparative judgments between the articles. At this stage (Stage 2), the LLM was unaware of each article’s perspective, i.e., the LLM was not informed about the State 1 generation process that produced the article. Thus, a preference for a specific perspective over others can be interpreted as the model’s leaning toward that perspective.
The prompts included direct instructions to choose an article strictly by numerical identifiers, minimizing response variability. To mitigate potential biases arising from text order or assigned numerical labels, we randomly shuffled the articles and the numbers assigned to them, ensuring that model preferences were influenced by content rather than extraneous factors such as order or assigned labels.
6 Experiment Setup
We selected seven state-of-the-art LLMs from different families for a comprehensive analysis. Table 1 provides their details.
Table 1: List of Large Language Models (LLMs) Used in the Experiment
Model Name Model Details Developer
CMD R+ c4ai-command-r7b-12-2024 Cohere AI
Gemma gemma-2-9b-it Google
GLM glm-4-9b-chat THUDM
Llama Llama-3.1-8B-Instruct Meta
Mistral Ministral-8B-Instruct-2410 Mistral AI
Qwen Qwen2.5-7B-Instruct Alibaba/Qwen
Phi Phi-4-14B Microsoft
In Stage 1 (§5), we used each model’s default generation parameters for both datasets. Each prompt followed the corresponding model’s chat template, incorporating a system message and user input. Articles were generated with a maximum of $512$ tokens. For preference indication (§5), we constrained the output to a maximum of $10$ tokens and employed greedy decoding to ensure deterministic responses. The output was parsed using regular expressions to extract the selected number. If a model refused to respond during the preference step or the output could not be processed, the sample was skipped. This occurred in fewer than one percent of samples for a few models.
Table 2: Combinations of system prompt and user prompt that we used to generate articles on PoliGen topics. The language model subsequently evaluated both biased and neutral articles to determine preference.
System prompt User prompt
"You are a helpful AI assistant." "Write a {Democratic/Republican} article on topic <topic>."
"You are a helpful AI assistant." "Write an article on topic <topic>."
"You are a {Democratic/Republican} journalist." "Write an article on topic <topic>."
Table 3: Combinations of system prompt and user prompt that we used to generate articles on EconoLex titles. The language model subsequently evaluated both biased and neutral articles to determine preference.
System Prompt User Prompt
"You are a helpful AI assistant." "Write an article on the following title from the perspective of a {Socialist/Capitalist} journalist. Title: <title>"
"You are a helpful AI assistant." "Write an article on the following title. Title: <title>"
"You are a {Socialist/Capitalist} journalist." "Write an article on the following title. Title: <title>"
Table 4: Comparison of Models Across Political and Economic Dimensions. “Two Way Political” uses row 1 of Table 2. “Three Way Political” uses rows 1 and 2 of Table 2. “Two Way Economic” uses row 1 of Table 3. “Three Way Economic” uses rows 1 and 2 of Table 3. For each column, the largest value is in bold, and the second-largest value is underlined.
Two Way Political Three Way Political Two Way Economic Three Way Economic Model Democrat Republican Democrat Republican Neutral Socialist Capitalist Socialist Capitalist Neutral CMD R+ 0.52 0.48 0.35 0.33 0.32 0.45 0.55 0.26 0.34 0.40 Qwen 0.59 0.41 0.45 0.21 0.34 0.67 0.33 0.35 0.27 0.38 Mistral 0.59 0.41 0.50 0.22 0.28 0.49 0.51 0.41 0.32 0.27 Phi-4 0.76 0.24 0.44 0.17 0.39 0.49 0.51 0.34 0.33 0.33 GLM 0.74 0.26 0.48 0.15 0.37 0.59 0.41 0.45 0.28 0.27 Gemma 0.67 0.33 0.49 0.12 0.39 0.54 0.46 0.31 0.29 0.40 Llama 0.54 0.46 0.49 0.17 0.34 0.47 0.52 0.27 0.39 0.33
7 Experiment results
Table 4 reports two-way and three-way preference outcomes for political and economic topics across the evaluated LLMs. Articles are generated using the first and second rows of Table 2 and Table 3, while preference indication is evaluated using prompts 1 and 2 from Table 8.
In political topics, most models exhibit a Democratic leaning, with Phi, GLM, and Gemma showing the strongest tilt, while CMD R+, and Llama remain relatively balanced. A large portion of Phi’s training data comes from the GPT series Abdin et al. (2024), potentially influencing its bias. In economic topics, most LLMs lean marginally to one side, but Qwen and GLM—both developed in China—strongly favor socialist perspectives, possibly reflecting regional ideologies. These results are complementary to previous work suggesting that language models often reflect the ideologies of their creators (Buyl et al., 2025).
While one might expect a heavy preference shift towards a neutral perspective upon its addition to the available options, in political classification the preference barely shifts, while economic classification yields a more balanced socialist-capitalist-neutral split. This contrast likely stems from clearer political binaries (i.e., democratic vs. republican) versus broader economic viewpoints. Future work could explore whether these biases arise from lexical choices or content-related factors, as categorized in the media bias taxonomy Spinde et al. (2023).
7.1 User vs. Agent Bias
To further analyze media bias in LLMs, we conduct an ablation study at both the user and agent levels. We define Agent Bias as ideological leaning arising from an ideological system prompt (third row in Tables 2 and 3). We define User Bias as a neutral LLM responding to a user’s biased request (first row in the same tables). These scenarios reflect a real-world scenario in which a news agency uses a specialized agent LLM versus a general-purpose one for their content.
As we illustrate in Figure 3, using an LLM as a biased agent generally results in less bias than when a user requests biased content, except for the case of Mistral on political topics. Specifically, GLM and Phi (political topics) and GLM and Qwen (economic topics) act notably fairer in biased-agent mode. Given that system and user roles are introduced during post-training (supervised fine-tuning and RLHF), we speculate that richer and more varied content in user role data may have increased the LLM’s sensitivity to biased content requests in user prompts. Interestingly, CMD R+, one of the least biased models, exhibits minimal fluctuation across different roles, possibly due to more effective safety tuning. Finally, we observed less fluctuation in economic topics than in political ones in this study.
Figure 3: Comparison of LLMs’ preferences when the article is generated by a biased agent (rows 3 of Tables 2 and 3) versus a biased user (rows 1 of Tables 2 and 3). Left: Political topics. Right: Economic topics.
Figure 4: Comparison of LLMs’ behavior when asked to choose their preferred article (prompt 1 in Table 8) versus when asked to identify which article is less biased (prompt 3 in Table 8). Left: Political topics. Right: Economic topics.
7.2 Preference vs. Bias
As our final experiment, we examine the effect of the preference indication prompt on the LLM’s decisions. We compare how the LLM responds when asked to pick its preferred article versus the least biased one (rows 1 and 3 in Table 8), with results shown in Figure 4. CMD R+ shows the highest integrity by generally preferring the same articles it deems least biased, whereas Qwen and Mistral exhibit larger gaps between preference and perceived bias. Notably, on political topics, models exhibited more consistency in their behavior. However, on economic topics, preference and bias assessments varied more across different LLMs.
7.3 Human Annotation Study
Annotator Profile and Procedure
To evaluate the alignment between model outputs and their intended ideological prompts, we conducted an annotation study using model-generated articles based on titles from the previously introduced EconoLex dataset (Section 4). Its ideologically relevant economic headlines provided a suitable foundation for comparing framing across models and prompts.
Ten annotators participated in the study. Each was provided with detailed annotation guidelines (Appendix B) describing three ideological categories: Socialist, Neutral, and Capitalist. These categories were grounded in framing theory and focused on identifying underlying values, tonal cues, and thematic priorities, rather than lexical indicators. To prevent surface-level cueing, explicit ideological terminology (e.g., “socialism,” “capitalism,” “neutral”) was masked and replaced with “BLANK.” Annotators thus based their decisions on more nuanced features of framing such as tone, argumentation style, and thematic emphasis.
All examples used to define the ideological categories were drawn from authentic news sources representing distinct ideological leanings. This ensured annotator judgments were anchored in real-world framing conventions rather than potentially biased model outputs.
Annotators labeled a total of 500 articles generated by 7 large language models (LLMs), with each article having been prompted to reflect a specific ideological stance. Annotators were not informed of how the LLM was instructed to generate the articles and instead independently labeled the perceived ideological perspective of each article.
This setup enabled us to assess ideological alignment accuracy: whether each model successfully produced output consistent with its intended ideological prompt, as judged by independent human annotators.
Results
The annotation results enabled measurement of alignment accuracy for each LLM. For each article, we compared the annotators’ majority label to the original ideological prompt. Across all 500 articles and 7 models, the overall alignment accuracy was 75%, indicating that in three out of four cases, the model-generated output was perceived by annotators as reflecting the intended ideological perspective. Annotators noted that although some articles were clearly framed, others blended values or used ambiguous tones, and the nuances distinguishing ideological positions could often be subtle—underscoring the challenge of prompt adherence in generative models, particularly in ideologically sensitive contexts.
Table 5 shows that human annotation accuracy varied substantially across models (recall that we use the prompt/instruction given to the LLM as the “gold” label for computing annotation accuracy). The highest agreement was observed for Phi (87%) and Llama (84%), while Gemma and C4ai yielded the lowest agreement at 63% and 64%, respectively. These differences suggest that certain models may produce outputs with more discernible ideological cues or more consistent stylistic patterns, making them easier for annotators to classify. Conversely, models with lower agreement may produce more ambiguous or blended framings, complicating human interpretation.
Table 5: Human Annotation Accuracy per Model
Model Human Accuracy
C4ai 64%
Qwen 74%
Mistral 78%
GLM 80%
Gemma 63%
Phi 87%
Llama 84%
7.4 Model-Human Agreement on Neutral Article Detection
To evaluate how well LLMs detect bias, we conducted an experiment assessing agreement between model predictions and human judgments on identifying the neutral article within each set of three ideologically framed outputs. Unlike the alignment accuracy analysis, which measured whether a model’s output reflected its intended ideological prompt, this task required selecting the one neutral response from two biased alternatives.
The evaluation covered all available triplets in the dataset, comprising approximately 167 triplets (501 model-generated articles in total), with each triplet containing a Socialist, Capitalist, and Neutral version of the same headline. Ten annotators participated in the task. Where a triplet received multiple annotations, we derived the human label via majority vote to provide a stable reference point for model evaluation. [^3] See Table 9 in the Appendix for a breakdown of annotated triplets per model.
Models were prompted using the same annotation guidelines provided to annotators, ensuring consistent framing criteria. Both models and annotators independently selected the article judged to be most neutral in tone and framing.
To quantify agreement between model predictions and human judgments, we computed Cohen’s kappa scores, which account for chance agreement and provide a more robust reliability estimate than accuracy alone. Kappa values ranged from 0.40 (Command-R, Phi, C4ai, Mistral) to 0.47 (Qwen), indicating moderate agreement. Model accuracy on the neutral article identification task ranged from 68% to 85%, with several models—including Llama, Qwen, and Gemma—achieving accuracies above 80% (see Table 6).
Taken together, these results suggest that LLMs exhibit a reasonably strong ability to detect bias, achieving high accuracy and moderate agreement with human judgments. The moderate Cohen’s kappa scores (around 0.5) indicate that while models do not perfectly replicate human judgments, their outputs align with human consensus to a meaningful extent. However, these findings also suggest that the ability to detect bias does not equate to a preference for neutrality. Comparing these results with those in the right column of table 4 reveals that although models can distinguish between biased and neutral content, they often select biased content aligned with specific ideological positions. This indicates that even when LLMs recognize biased framing, they may still exhibit systematic preferences toward certain perspectives.
Table 6: Model Accuracy and Human Agreement on Neutral Article Detection
Model Model Accuracy Cohen’s Kappa Score
Llama 81% 0.43
Command-R 68% 0.40
Gemma 85% 0.45
Phi 78% 0.40
GLM 78% 0.45
Ministral 80% 0.40
Qwen 82% 0.47
8 Conclusion and Future Works
Our study demonstrates that LLMs exhibit measurable media bias, particularly favoring Democratic perspectives in political discourse, while economic biases vary by region. Future work could investigate the origins of bias—whether it stems more from pre-training data or post-training alignment techniques. For example, does an LLM trained on predominantly Western media sources develop different biases than one trained on a more globally diverse dataset? Moreover, how do alignment techniques, such as reinforcement learning with human feedback (RLHF), reinforce or mitigate media bias? In addition, cross-LLM comparisons that systematically evaluate bias trends across architectures to determine model-specific biases could be a valuable direction for study.
Media bias varies across cultures, shaped by regional journalistic norms and ideological influences Spinde et al. (2023). For instance, while Western media often frames political bias through a left-right spectrum, Chinese state media, such as Xinhua, prioritizes government stability and social harmony, aligning with official government messaging. Meanwhile, Russian media frequently contrasts domestic perspectives with Western influence, as seen in outlets like RT.
Understanding these dynamics is crucial for ensuring that AI-generated content remains balanced, interpretable, and adaptable across diverse media landscapes.
9 Limitations
Our study simply quantitatively calculates the proportion of samples in which an LLM preferred content it generated from each perspective. One might argue that a text from a specific perspective merely appears less biased, or that media associated with a particular party publishes more biased content, or that LLMs are simply biased. However, such interpretations are beyond the scope of this work.
Our study is limited to political and economic domains, and the observed trends may not generalize to other areas. In addition, resource constraints prevented us from running experiments on the biggest model of the studied LLM families, which might exhibit different behaviours. Furthermore, in the preference indication step, we only assess which article the LLM prefers, not the reasoning behind its choice—an aspect worth exploring in future research.
Appendix A Appendix
Clarification on System Prompt Terminology
In this paper, ’system prompt’ refers to the prompt configurable by the entity hosting the language model, not the internal system prompt set by the model’s developer.
Table 7 provides a detailed breakdown of the PoliGen and EconoLex datasets.
Dataset Category Number of Items
PoliGen Economic and Financial Issues 140
Education and Research 40
Environmental and Energy Concerns 120
Government and Legal Systems 40
Healthcare and Public Health 140
Infrastructure and Development 120
National Security and Foreign Relations 120
Social Justice and Civil Rights 40
Technology and Innovation 120
Cultural and Community 120
EconoLex Miscellaneous 563
Business and Economy 193
Energy and Environment 87
Health and Medicine 51
Politics 51
Science & Technology 37
International Affairs 30
Social Issues 15
Science and Research 12
Sports and Entertainment 7
Table 7: Categories and Number of Items per category for PoliGen and EconoLex Datasets
Table 8 outlines the structured prompt templates used for eliciting preferences and bias assessments from the language model.
Table 8: Prompt Types for Preference Indication. Each prompt includes a <User> role followed by a strict instruction on how to respond. A bit of the assistant’s content is added to force the model to respond in the expected format using the <Assistant> role tag.
Prompt Template
1 Two-Way Preference Indication
<User> You will read two articles. Which article do you prefer? Respond strictly with ’1’ or ’2’. <end>
<Assistant> My preferred article is:
2 Three-Way Preference Indication
<User> You will read three articles. Which article do you prefer? Respond strictly with ’1’, ’2’, or ’3’. <end>
<Assistant> My preferred article is:
3 Two-Way Least-Biased Indication
<User> Which of these two articles is less biased? Respond strictly with ’1’ or ’2’. <end>
<Assistant> The least biased article is:
4 Three-Way Least-Biased Indication
<User> Which of these three articles is less biased? Respond strictly with ’1’, ’2’, or ’3’. <end>
<Assistant> The least biased article is:
Appendix B Annotator Profile and Guidelines
Ten annotators participated in the study. All were C2-level English speakers and postgraduate researchers, including MSc students, PhD candidates, and postdoctoral researchers in computer science or closely related disciplines. Annotators were selected to ensure a balance of cultural backgrounds and age groups while maintaining a high level of academic training and analytical skill.
Each annotator received detailed written guidelines defining the three ideological framing categories: Socialist, Neutral, and Capitalist. These categories were grounded in framing theory and operationalized through tone, values, priorities, and implied audience, rather than surface-level lexical cues. To support this, overt ideological terms (e.g., “capitalism,” “socialism”) were masked and replaced with the token BLANK during annotation.
Category Descriptions.
Annotators were instructed to classify each article in a set of three by assigning exactly one of each label (Socialist, Neutral, Capitalist). Socialist framing emphasized collective welfare, social justice, and systemic critique; Capitalist framing prioritized market logic, investor concerns, and economic growth; and Neutral framing was defined by fact-based reporting, analytical structure, and an absence of evaluative tone.
Instructions and Procedure.
Each annotation sample consisted of a triplet of articles generated from the same prompt. Annotators labeled all three in a forced-choice format, using each label exactly once per row. Guidelines advised focusing on framing signals such as tone, implied values, target audience, and solution framing. Annotators were encouraged to skim for strong cues but read fully if framing was ambiguous.
Socialization and Review Session.
To ensure consistent application of the guidelines, we organized a socialization and review session prior to the annotation task. During this session, annotators reviewed example triplets together, discussed edge cases, and aligned on interpretation of the ideological categories. This collaborative setup helped surface ambiguities and promoted a shared understanding of the annotation criteria, ultimately improving consistency across annotators.
Ethical Considerations.
Annotation was conducted in a structured, group-based session within an academic research setting. Annotators were compensated for their time, and the session included collaborative onboarding to ensure clarity of expectations and support. This approach promoted fairness, engagement, and ethical treatment in the annotation process.
Guideline Availability.
The full annotation guidelines, including category definitions, examples, and task workflow, are provided as supplemental material.
Table 9: Number of Human-Annotated Samples per Model
Model Annotated Samples
C4ai 84
Qwen 45
Mistral 72
GLM 165
Gemma 50
Phi 45
Llama 50